A dielectric of dielectric constant 5.0 is filled in the gap between the plates of a capacitor.calculate the factor by which the capacitance is increased,if the dielectric is only sufficient to fill up 1/5 of the gap.


1
Expert's answer
2021-01-25T15:59:39-0500

"C_0=\frac{\epsilon_0S}{d}"


"C=C_1+C_2=\frac{\epsilon\epsilon_0(1/5)S}{d}+\frac{\epsilon_0(4/5)S}{d}="


"=\frac{\epsilon_0S}{d}(\epsilon/5+4/5)" .


"k=\frac{C}{C_0}=\epsilon/5+4/5=5/5+4/5=1.8" .Answer